OpenStudy (anonymous):

factor out 6g^2+11-35 step by step pleasee!!

OpenStudy (anonymous):

6g^2+11-35 = 6g^2 -24 =6*(g^2-4) =6(g+2)(g-2)

OpenStudy (anonymous):

where did you get the 24 in the equation?

OpenStudy (turingtest):

11-35=-24

OpenStudy (turingtest):

not supposed to be 6g^2+11g-35 is it?

OpenStudy (anonymous):

okay, the equation is 6g^2+11g-35 okay i know the first step is to times 6 and 35 and you get -210 and just need 2 numbers that when multiplied get to -210 but added 11

OpenStudy (anonymous):

please help!!

OpenStudy (anonymous):

6g^2+21g-10g-35 =3g(2g+7)-5(2g+7) =(3g-5)(2g+7)

OpenStudy (anonymous):

 thats were i get stuck..i need to numbers that multiplied equal -210 but added 11

OpenStudy (anonymous):

ksram you are my HERO!!!!!!